OpenStudy (anonymous):

George weighs 122 pounds. Together, he and his sister weigh four pounds more than three times the weight of his sister. What is the weight, w, of George’s sister?

OpenStudy (anonymous):

x = weight of his sister; x+122 = 3x +4; 2x = 118 => x= 59;

OpenStudy (anonymous):

you asume that his sister is x; what the question say that , George + his_sister = 3*his_sister +4